Ria sees a bird outside her bedroom window. She says the bird moved 40 m East, 30 m North and 60 m west before it disappeard. What was the displacement of the bird while Ria was watching it?
1
Expert's answer
2021-10-13T16:56:40-0400

First, find displacement in both vertical (northward) and horizontal (eastward) directions:


"D_N=30\text{ m},\\
D_E=40-60=-20\text{ m}."

Second, find the magnitude of the resultant with Pythagorean theorem:


"D=\sqrt{D_N^2+D_E^2}=36\text{ m}."